OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to the line 2x -3y = 12 and passes through the point (2, 6).

OpenStudy (anonymous):

perpendicular of y=2/3x-4 is y=-3/2x, then use the equation for finding intercept, (y-y1)=-3/2(x-x1)+b to solve

OpenStudy (anonymous):

? can you help me like step by step

OpenStudy (p0sitr0n):

\[ m2 = -1/m1\] this is perpendicular slope replace in the equation. gives y = -3/2x +b replace point cooordinates to find b gives 6 = -3/2 * 2 +b b =9 y = -3/2x+9

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

is that my answer?

OpenStudy (p0sitr0n):

we only know that the slope is prependicular. it is the opposed of the inverse of another slope , by definition next, when we have the m2, we replace a known point couple to determine an unknown value b the following is how to know the slope 2 x -3y =12 -3y = 12 -2x y = -4+2/3x the slope is before the x, so it is 2/3 applay formula, you get - 1 / 2/ 3 = -3/2